OpenStudy (anonymous):

Solve the equation for x 0<= x <=2pi sin^2x - sqrt(2)cosx = cos^2x + sqrt(2)cosx + 2

OpenStudy (anonymous):

i got 0=cos^2x + sqrt(2)cosx + 1 I tried using the quadratic formula but i get a discriminate with a negative.

OpenStudy (anonymous):

help?

OpenStudy (mertsj):

\[\sin ^2x-\sqrt{2}\cos x=\cos ^2x+\sqrt{2}\cos x+1\]

OpenStudy (mertsj):

Is that the problem or is there parentheses around x+1 ?

OpenStudy (anonymous):

yes, tht is the problem except it is + 2 not + 1

OpenStudy (anonymous):

\[\sin^2x - \sqrt{2}cosx = \cos^2x + \sqrt{2}cosx + 2\]

OpenStudy (mathstudent55):

\(\color{red}{\sin^2x} - \sqrt2 \cos x = \cos^2x + \sqrt2 \cos x + 2\) Use the identity \(\sin^2 x = 1 - \cos^2 x\) \( \color{red}{ 1 - \cos^2 x } - \sqrt2 \cos x = \cos^2x + \sqrt2 \cos x + 2\) \(2 \cos^2 x + 2 \sqrt{2} \cos x + 1 = 0 \)

OpenStudy (mertsj):

Now use the quadratic formula.

OpenStudy (mathstudent55):

Use the quadratic formula to solve for \( \cos x\) and then use trig to solve for x.

OpenStudy (anonymous):

wow, thank you guys so much!

OpenStudy (mathstudent55):

wlcm